OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 4^y=5

OpenStudy (anonymous):

I can't post the graphs... can somebody please tell me how to graph this?

OpenStudy (anonymous):

It's logarithmic, if that helps at all.

OpenStudy (solomonzelman):

\(\large\color{black}{ 4^y=5}\) \(\large\color{black}{ \log 4^y=\log 5}\) \(\large\color{black}{ y~\log 4=\log 5}\) \(\large\color{black}{ y=\log 5 \div \log 4}\) \(\large\color{black}{ y=\log_45}\)

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/gk8bjtlxkg

OpenStudy (solomonzelman):

Just a vertical line

OpenStudy (anonymous):

Uh oh... All of my graphs have curves...

OpenStudy (anonymous):

A curve. excuse me.

OpenStudy (solomonzelman):

\(\large\color{black}{ \log_45=1.16}\)

OpenStudy (solomonzelman):

Just a vertical line

OpenStudy (solomonzelman):

maybe you mean \(\large\color{black}{ 4^y=x}\) ?

OpenStudy (anonymous):

Unfortunately, no. I typed it how the problem did.

OpenStudy (solomonzelman):

well, y is in the exponent, but in this case, you have x set to zero, and all you get is the approximation of y=1.16. it is just a vertical line, as far as I know...

OpenStudy (solomonzelman):

http://www.wolframalpha.com/input/?i=4%5Ey%3D5

OpenStudy (anonymous):

Ok. Maybe I'm just missing something... oi

OpenStudy (solomonzelman):

the blue is the exponential that you are talking about

OpenStudy (anonymous):

OK- Unfortunately, none of my graphs look like that.

OpenStudy (anonymous):

I've got it! The 1.16 is an asymptote, so I just need to look for a graph that looks like it has an asymptote just above 1!

OpenStudy (anonymous):

Yay!

OpenStudy (solomonzelman):

yes, apparently it is so...

OpenStudy (anonymous):

Thank you!